OpenStudy (anonymous):

HOW TO GET THE ANSWER?? The function f(x) = 582(7)x represents the growth of a mosquito population every year in a remote swamp. Troy wants to manipulate the formula to an equivalent form that calculates every 2 months, not every year. Which function is correct for Troy's purposes? f(x) = 582(7 to the one sixth power)6x f(x) = 52(7)x f(x) = 582(712)the x over 12 power f(x) = 5852(7)x

OpenStudy (anonymous):

HELP!

OpenStudy (kirbykirby):

oh is that \(\large f(x)=5852(7)^x\)

OpenStudy (anonymous):

d?

OpenStudy (kirbykirby):

no sorry I was just confused about your notation

OpenStudy (anonymous):

yeah but you mean 582*

OpenStudy (anonymous):

is it c?

OpenStudy (kirbykirby):

It looks like a) should be right but I'm not sure of the formatting of the answer. But really if : \[x = \text{1 year}\] then \[\frac{x}{12}=\text{1 month}\], so \[\frac{2x}{12}=\frac{x}{6}=\text{2 months} \]

OpenStudy (anonymous):

So just go with a?

OpenStudy (kirbykirby):

i think so

OpenStudy (anonymous):

thank you!

OpenStudy (kirbykirby):

:)

OpenStudy (whpalmer4):

\[f(x)=582*7^x = 582*(7^{1/6})^{6x}\]because \[(a^b)^c = a^{b*c}\] Following this pattern, if we wanted a formula that worked on a daily basis, we would use \[f(x) = 582*(7^{1/365})^{365x}\] And so on...